In a perfect competitive market the price of the commodity is 3 per unit of total output.the total cost function of a firm is given c =(1/3) - 4q2+ 10 q + 10.

Find the profit maximization output and total profit


1
Expert's answer
2021-01-13T10:23:01-0500
"TC=\frac{1}{3}-4q^2+10q+10"


"MC=-8q+10"


"TR=3q"


"MR=3"


"MR=MC"


"-8q+10=3"


"q=0.875"


"\pi=TR-TC=4q^2-7q-10\frac{1}{3}"

"\pi=-13.4"